OpenStudy (anonymous):

Solve: 6(x + 4) = 8(x − 2) −4 4 −20 20

OpenStudy (anonymous):

20

OpenStudy (anonymous):

do you need an explination?

OpenStudy (anonymous):

6(x+4)=8(x-2) = 6x + 24 = 8x - 16 8x - 6x = 24 + 16 2x = 40 x = 20

OpenStudy (anonymous):

alright ty

OpenStudy (anonymous):

np